The star Tau Ceti's mass is 1.6 ✕ 1030 kg, its radius is 5.5 ✕ 105 km, and it has a rotational period of approximately 34 days. If Tau Ceti should collapse into a white dwarf of radius 7.5 ✕ 103 km, what would its period (in s) be if no mass were ejected and a sphere of uniform density can model Tau Ceti both before and after?
Answer #1

f) w = Kf-ki = 11.639- 2.7597 = 8.28 kJ M=16 x 1030 Kg, R. = 5.5x105 km = 5.5*108m R2 = 7.5x 10% km = 4.5 x106m, T, = 34 days